OpenStudy (anonymous):

A piggy bank contains $2.40 in nickels and dimes. If there are 33 coins in all, how many nickels are there?

OpenStudy (blacksteel):

Let's let the number of nickels be x. Then the number of dimes must be 33-x. Then 5x + 10(33-x) = 240 5x + 330 - 10x = 240 330 - 5x = 240 5x = 90 x = 18